# Accelerating Quantum Eigensolver Algorithms With Machine Learning

### v1

@@ 5.1 Data Augmentation
The accuracy and generalisation of an ML model are closely linked to the size of the dataset. In the context of quantum simulation data, each data instance encompasses numerous features due to the substantial memory required to represent the Hamiltonian. To generate a sufficiently large dataset for training, we utilise a classical state vector simulator rather than a quantum device, as quantum computing time is prohibitively expensive. This simulator allows us to compute the exact energy levels of smaller Hamiltonians (i.e. fewer than 28 qubits) with high precision. Each data instance in our training dataset encompasses the complete set of information required for the quantum problem, including the Hamiltonian [^4] In practice, we do not include in our data the Hamiltonian as-is but a compressed representation of the Hamiltonian to save memory and avoid overfitting. Compressing a Hamiltonian in the context of QC reduces the complexity of the Hamiltonian operator but keeps the essential physical properties of the system. We compressed by truncating small terms., its exact energy level, and the hyperparameters used to compute this level.
@@
Note on the energy levels predictions: As energy levels are real numbers (i.e., are in a continuous domain), classification might result in precision issues during the prediction phase. Moreover, the exact energy level value is irrelevant as we do not use these values directly. We are interested in knowing whether certain data (Hamiltonian and some quantum-system-related parameters) lead to further minimised answers (the lowest energy level value). This is the relationship between these values. Therefore, we have used a regressor to predict roughly the lowest energy levels.
Figure 3: Classical pre-processing phase - training a regressor on smaller systems
@@ 5.2 AccelerQ
To ensure the generalisability of AccelerQ and its applicability to various QE algorithms (e.g., ADAPT-QSCI or QCELS), we have developed a versatile wrapper. This wrapper encompasses the Hamiltonian, the number of qubits, a parameter indicating whether the computation should be performed classically or on a quantum computer, and the hyperparameters (of both: the QE and ML problems, while we only optimise here those of the QE problem). In our evaluation (see Section 6), we applied AccelerQ to the ADAPT-QSCI and the QCELS algorithms.
@@
We add open-source commonly used molecular Hamiltonians (H2O, LiH, BeH2, Hemocyanin [2], and H chain) and combine them with the ones provided by the challenge Hamiltonians of $16$ qubits or less, details in [7]. The wrapper is run in classical mode while giving a randomly generated set of hyperparameters each time, recording the energy level as the score. This process is repeated to produce a data array for training a Regularising Gradient Boosting Regression (XGBoost). The generated data—consisting of a compressed Hamiltonian and hyperparameter vectors (Xs) and their corresponding energy levels (Ys)—is stored for further processing. This stage corresponds to the first of the three steps depicted in Figure 3. We utilise a classical eigensolver during data preparation (Subsection 5.1).
@@ 5.3 Implementation Details
Our implementation is divided into three parts (data collection, training and optimisation) and is written in Python 3.10.12. The rest of the requirements are derived from the rules of the 2024 Quantum Algorithm Grand Challenge (QAGC2024), with one exception: we utilised GPU during the training phase of our models while keeping them relatively small to be able to deploy them on desktop CPU. We used the Python library QURI Parts [^5] https://pypi.org/project/quri-parts/ to simulate and access QC.
@@
Participants in the 2024 Quantum Algorithm Grand Challenge (QAGC2024) were tasked with challenging the ADAPT-QSCI Algorithm [40], a quantum-classical hybrid approach designed to compute the ground states and energies of many-body quantum Hamiltonians (Subsection 3.1). The challenge specifically focuses on optimising a quantum algorithm for solving a $28$-qubit system under several constraints (see full list [47]). These constraints included a specific machine specification (CPU, 16 GB RAM), a practical time limit of $6\times 10^{5}$ seconds to ensure all solutions could be evaluated within a reasonable timeframe and a shots limit of $10^{7}$ due to the high cost of quantum computer execution. Participants were required to make essential use of quantum computers, as the goal was a quantum competition, and current classical algorithms do not scale efficiently for systems with $28$ or more qubits.
Furthermore, participants were instructed to output only the energy level of the ground state, calculated using the quantum algorithm of choice. Exact calculations of the ground-state energy using methods like the Bethe ansatz were prohibited, as such methods do not scale well for larger systems ($28$ qubits or more). Additionally, participants were not allowed to hard-code initial parameters or hyperparameters to ensure adaptability to new Hamiltonians. That constraint ensured that the prosposed solution is generic and not tailored for a specific Hamiltonian or subset of problems.
To facilitate the automatic evaluation of all participants’ solutions, the contest organisers provided a specific structure for their responses, limiting participants to modifying a single answer file. While this constraint streamlined the evaluation process, it also introduced certain limitations. Notably, it prevented us from leveraging the full potential of machine learning techniques, as models could not be stored globally and had to be retrained for each evaluation. Although this approach was effective for the contest evaluation, it is not ideal for software engineering practices that prioritise modularity. A more effective strategy would be to divide the problem into distinct stages, each with its own constraints. Constraints provided by the contest organisers are particularly relevant for the prediction phase but may not be suitable for the training phase. Recognising these limitations, in future iterations of our solution we plan to adopt further a more flexible approach, allowing for greater use of advanced methods and better adaptability, while respecting the original framework and intentions set by the challenge organisers.
@@ 6 Evaluation
We evaluated AccelerQ for its ability to further optimise ADAPT-QSCI and QCELS algorithms by suggesting better hyperparameters tailored per system.
6.1 Methodolgy
We demonstrated our idea on the optimisation of hyperparameters of two quantum eigensolver algorithms. We used 4- to 16-qubit systems to train the models. We deployed the models on 20-, 24- and 28-qubit systems. We aim to understand:
RQ1: To what extent can AccelerQ’s optimisation of hyperparameters alone accelerate and improve the efficiency and accuracy of Quantum Eigensolver (QE) algorithms on NISQ devices in terms of runtime, error, and system size?
@@ 6.1 Methodolgy
In our evaluation, we applied our methodology to 16 larger systems of 20-, 24- and 28-qubit with known lowest energy levels, using the two QE implementations: ADAPT-QSCI and QCELS (Section 4) to answer RQ1 and RQ2.
We constructed two models—one for each implementation—using data extracted classically (Section 4) from up to 16-qubit systems. The data extraction and model training were general. The resultant model aimed to predict the optimal hyperparameters for its QE implementation and a system (Hamiltonian). The optimisation process took an implementation, its trained model and a system (a problem to solve) as input and returned the optimal hyperparameters for this setup.
@@
Our evaluation baseline is the results obtained with an implementation’s default hyperparameters, fixed across all systems in the evaluation. The hyperparameters were specific per implementation, were optimised, and their performance was compared against the baseline. Apart from the hyperparameters, ADAPT-QSCI and QCELS implementations were provided with ham and number_qubits, the input system and the required number of qubits for the corresponding Hamiltonian, and the flag is_classical being set to True during the data collection phase, and either True or False otherwise (classical or QC). We did not optimise this part.
@@
ADAPT-QSCI. The ADAPT-QSCI implementation [40] includes several hyperparameters that control its operation. These hyperparameters have default values, which we used as our baseline for comparison. num_pickup (default: 100) and coeff_cutoff (default: 0.001) parameters control the terms retained or removed from the Hamiltonian compressed representation. self_selection indicates whether self-selection is enabled (default: False), which forces the algorithm to work in subspace. iter_max is the maximum total number of iterations (default: 100). sampling_shots is the number of sampling shots for measurements per iteration (default: $10^{5}$). atol is the absolute tolerance for convergence criteria (default: $61e-$ 6). final_sampling_shots_coeff is the number of shots for the calculations if the same operator appears twice or the operator parameter is close to zero (default: 5). num_precise_gradient is the number of operators from the pool to calculate gradient more precisely (default: 128). max_num_converged specifies the maximum number of iterations within atol needed for a solution to be considered converged (default: 2). reset_ignored_inx_mode specifies after how many iterations previously used operators can be reused (default: 0). These default values are the baseline for configuring the ADAPT-QSCI implementation, but these can be overridden (e.g.) when using AccelerQ suggestions tailored per system.
@@
QCELS. The QCELS hyperparameters are a different set of arguments. This includes the following: ham_terms is the number of individual terms that are retained in the Hamiltonian after truncation. The Hamiltonian is transformed to a qubit Hamiltonian formed from a linear combination of Pauli strings, the truncation then retains the largest ham_terms of these and the rest are discarded. The default value is set to be 200 and we search for its optimal value between random.randint(50, 1000). ham_cutoff is the same as coeff_cutoff in ADAPT-QSCI implementation, setting a minimum value for the retained coefficients, discarding all terms in the Hamiltonian with coefficients lower than this threshold. delta_t is the time step for the simulation or evolution of the system (the default is 0.03; delta_t = random.uniform(1e-3, 0.3)), the expectation value of the time evolution operator is calculated at a set of times starting from zero and separated by this value. n_Z is the total number of points used in fitting the time evolution and so n_Z11-1- 1 is the total number of points at which the expectation value is evaluated. The default value is 10; n_Z = random.randint(5, 25). alpha is a scalar parameter that determines how much smaller the parameters $r_{2}^{(2)}$, $\theta_{2}^{(2)}$ and $r_{2}^{(3)}$, $\theta_{2}^{(3)}$, $\theta_{3}^{(3)}$ are forced to be than (default is $0.8$ ;alpha = random.uniform(0.5, 1)).
@@ 6.3 Results
Table 1: Predicted optimal hyperparameters (A-J col.) with the first row stating the default values for ADAPT-QSCI implementation (Size: the approximate size of the Hamiltonian as a vector of terms, A: num_pickup, B: coeff_cutoff, C: self_selection, D: iter_max, E: sampling_shots, F: atol, G: final_sampling_shots_coeff, H: num_precise_gradient, I: max_num_converged, and J: reset_ignored_inx_mode)
System Size A B C D E F G H I J
Default - 100 0.001 0 100 100000 1.00E-06 5 128 2 0
20qubits_00 63636 985 7.64647e-03 0 344658 49458 5.67168e-05 3 77 2 1
20qubits_01 54717 807 9.41494e-03 1 382109 964853 9.00118e-05 6 79 2 64
20qubits_02 54723 934 4.56117e-03 0 478268 875460 4.27314e-05 8 294 3 83
20qubits_03 54710 551 1.13139e-03 1 802530 106374 7.23056e-05 1 161 2 79
20qubits_04 63636 182 4.06265e-03 0 398085 61950 7.69414e-05 7 194 4 11
20qubits_05 46 593 8.46933e-04 1 278431 148423 6.66628e-05 7 158 4 58
24qubits_05 56 67 7.77741e-03 1 655715 676443 3.37411e-05 8 176 2 89
24qubits_06 56 292 4.00805e-03 0 762863 603105 6.33079e-05 3 159 4 82
24qubits_07 56 80 4.06535e-03 0 659716 711486 9.45728e-05 4 37 3 0
24qubits_08 56 215 6.59156e-03 1 885413 291401 1.74522e-05 1 247 4 81
24qubits_09 56 716 7.29064e-03 0 401350 929361 6.82934e-05 5 257 2 65
28qubits_01 98700 1000 6.58567e-03 0 615787 300938 1.04770e-05 8 224 4 98
28qubits_00 98700 197 1.42217e-03 1 460054 209518 7.62522e-05 7 287 3 54
28qubits_02 98700 531 5.88032e-03 1 402352 258683 4.81220e-06 8 277 2 47
28qubits_03 98700 629 4.35568e-03 0 429007 82966 3.89304e-05 8 239 4 8
28qubits_04 98700 698 7.96341e-03 0 221061 633412 7.86532e-05 4 68 4 89
@@
Table 2: Predicted optimal hyperparameters (A-E cols.) with the first row stating the default values for QCELS implementation (Size: the approximate size of the Hamiltonian as a vector of terms, A: delta_t, B: n_Z, C: ham_terms, D: ham_cutoff, and E: alpha)
System Size A B C D E
Default - 0.03 10 200 1e-9 0.8
20qubits_00 63636 1.37929e-01 14 877 6.77275e-03 9.76144e-01
20qubits_01 54717 2.87833e-01 8 69 9.23310e-04 5.26429e-01
20qubits_02 54723 1.05366e-01 10 989 4.31098e-04 8.77018e-01
20qubits_03 54710 1.04235e-01 16 68 6.91101e-03 5.12876e-01
20qubits_04 63636 1.31645e-01 19 971 2.10235e-03 7.76523e-01
20qubits_05 46 1.98379e-01 7 187 8.34526e-03 5.33536e-01
24qubits_05 56 2.08699e-01 24 933 5.75035e-03 6.20185e-01
24qubits_06 56 1.75486e-01 10 398 7.28888e-03 9.72887e-01
24qubits_07 56 1.80166e-01 11 286 1.06348e-03 9.75924e-01
24qubits_08 56 2.72722e-01 22 453 6.67156e-03 8.29813e-01
24qubits_09 56 2.09069e-01 24 650 3.99377e-03 5.20032e-01
28qubits_00 98700 7.00174e-03 7 310 6.71154e-03 6.32763e-01
28qubits_01 98700 1.99713e-01 23 309 8.39957e-03 5.54053e-01
28qubits_02 98700 5.46760e-02 12 997 1.22706e-03 5.60243e-01
28qubits_03 98700 4.23353e-02 6 553 8.40543e-03 8.66162e-01
28qubits_04 98700 2.33005e-01 25 280 1.30071e-04 5.98526e-01
Table 3: Hamiltonian sizes (Size col.) and the size after compression used when querying the ML Models (Comp. col.). The column Comp. represents the size of the system as utilised in the ML model. We marked with an asterisk the open-source molecular Hamiltonians.
System Size Comp.
20qubits_00 63636 91
20qubits_01 54717 287
20qubits_02 54723 445
20qubits_03 54710 338
20qubits_04 63636 65
20qubits_05(*) 46 46
System Size Comp.
24qubits_05(*) 56 48
24qubits_06(*) 56 46
24qubits_07(*) 56 50
24qubits_08(*) 56 51
24qubits_09(*) 56 47
System Size Comp.
28qubits_00 98700 167
28qubits_01 98700 158
28qubits_02 98700 98
28qubits_03 98700 144
28qubits_04 98700 213
@@
Table 1 and Table 2 show the optimal hyperparameters found by AccelerQ with each Hamiltonian (System col.) employing the two models above. Size col. is the number of terms in the system. The first row shows the default values of the implementation. Table 1 shows the ADAPT-QSCI’s optimal hyperparameters, with col. A-J represents the predicted optimal hyperparameters values. Similarly, Table 2 presents the optimal hyperparameters for the QCELS implementation in col. A-E. The value presented in Table 1 for the predicted iter_max was capped during execution [^7] The platform automatically stopped the computation once the maximum number of shots, 10 000 0001000000010\,000\,00010 000 000, was reached. to ensure that we do not have unlimited resources. In practice, the iter_max has the same limit as with the default parameters (i.e. iter_max=1e7/sampling_shots; col. E and col. D). Table 3 shows the size (Size col.) and the compressed size (Comp. col.) of the Hamiltonians (System col.) used when querying the ML model during prediction, marking with an asterisk the open-source molecular Hamiltonians. Different cutoffs and compression rates were used for ML part than in the QE implementations. Consequentially, the sizes recorded in Table 3 are true only for the ML model.
In Table 3, the number of terms in the open-source molecular Hamiltonians systems ($\leq 46$) is significantly smaller than the challenge Hamiltonians ones ($\geq 54k$), even after compression, the open-source molecular Hamiltonians systems was still much smaller on average. This suggests that these systems may differ to some extent.

### v2

@@ 5.1 Data Augmentation
The accuracy and generalisation of an ML model are closely linked to the size of the dataset. In the context of quantum simulation data, each data instance encompasses numerous features due to the substantial memory required to represent the Hamiltonian. To generate a sufficiently large dataset for training, we utilise a classical state vector simulator rather than a quantum device, as quantum computing time is prohibitively expensive. This simulator allows us to compute the exact energy levels of smaller Hamiltonians (i.e. fewer than 28 qubits) with high precision. Each data instance in our training dataset encompasses the complete set of information required for the quantum problem, including the Hamiltonian [^4] In practice, we do not include in our data the Hamiltonian as-is but a compressed representation of the Hamiltonian to save memory and avoid overfitting. Compressing a Hamiltonian in the context of QC reduces the complexity of the Hamiltonian operator but keeps the essential physical properties of the system. We compressed by truncating the smallest terms, its exact energy level, and the hyperparameters used to compute this level.
@@
Note on the energy levels predictions: As energy levels are real numbers (i.e., are in a continuous domain), classification might result in precision issues during the prediction phase. Moreover, the exact energy level value is irrelevant as we do not use these values directly. We investigate whether specific data (Hamiltonian and quantum-system-related parameters) contribute to further minimising the lowest energy level. We are interested in exploring the relationship between these values. Therefore, we have used a regressor to predict roughly the lowest energy levels.
Figure 3: Classical pre-processing phase - training a regressor on smaller systems.
@@ 5.2 AccelerQ
To ensure the generalisability of AccelerQ and its applicability to various QE algorithms (e.g., ADAPT-QSCI or QCELS), we have developed a versatile wrapper. This wrapper encompasses the Hamiltonian, the number of qubits, a parameter indicating whether the computation should be performed classically or on the challenge MPS simulator, and the hyperparameters (of both, the QE and ML problems, while we only optimise here those of the QE problem). In our evaluation (see Section 6), we applied AccelerQ to the ADAPT-QSCI and the QCELS algorithms.
@@
We use Hamiltonians from [58] and open-source commonly used molecular Hamiltonians (H2O, LiH, BeH2, Hemocyanin [2], and Hydrogen chain) of $16$ qubits. The wrapper is run in classical mode while giving a randomly generated set of hyperparameters each time, recording the energy level as the score. This process is repeated to produce a data array for training a Regularising Gradient Boosting Regression (XGBoost). The generated data—consisting of a compressed Hamiltonian and hyperparameter vectors (Xs) and their corresponding energy levels (Ys)—is stored for further processing. This stage corresponds to the first of the three steps depicted in Figure 3. We utilised a quantum eigensolver run on an emulator during data preparation (Subsection 5.1).
@@ 5.3 Implementation Details
Our implementation is divided into three parts (data collection, training and optimisation) and is written in Python 3.10.12. The rest of the requirements are derived from the rules of the 2024 Quantum Algorithm Grand Challenge (QAGC2024), with one exception: we utilised GPU during the training phase of our models while keeping them relatively small to be able to deploy them on desktop CPU. We used the Python library QURI Parts [^5] https://pypi.org/project/quri-parts/ to simulate the QC.
@@
Participants in the 2024 Quantum Algorithm Grand Challenge (QAGC2024) were tasked with challenging the ADAPT-QSCI Algorithm [49], a quantum-classical hybrid approach designed to compute the ground states and energies of many-body quantum Hamiltonians (Subsection 3.1). The challenge specifically focuses on optimising a quantum algorithm for solving a $28$-qubit system under several constraints (see full list [59]). These constraints included a specific machine specification (CPU, 16 GB RAM), a practical time limit of $6\times 10^{5}$ seconds to ensure all solutions could be evaluated within a reasonable timeframe and a shots limit of $10^{7}$ due to the high cost of quantum computer execution. Participants were required to make essential use of quantum computers, as the goal was a quantum competition, and current classical algorithms do not scale efficiently for systems with many ($\gtrsim 50$) qubits.
Furthermore, participants were instructed to output only the energy level of the ground state, calculated using the quantum algorithm of choice. Exact calculations of the ground-state energy using methods like the Bethe ansatz were prohibited, as such methods are either inefficient or do not work for general systems. Additionally, participants were not allowed to hard-code initial parameters or hyperparameters to ensure adaptability to new Hamiltonians. That constraint ensures that the proposed solution is generic and not tailored for a specific Hamiltonian or subset of problems.
To facilitate the automatic evaluation of all participants’ solutions, the contest organisers provided a specific structure for their responses, limiting participants to modifying a single answer file. While this constraint streamlined the evaluation process, it also introduced certain limitations. Notably, it prevented us from leveraging the full potential of machine learning techniques, as models could not be stored globally and had to be retrained for each evaluation. Although this approach was effective for the contest evaluation, it is not ideal for software engineering practices that prioritise modularity. A more effective strategy would be to divide the problem into distinct stages, each with its own constraints. Constraints provided by the contest organisers are particularly relevant for the prediction phase but may not be suitable for the training phase. Recognising these limitations, in future iterations of our solution, we plan to adopt a more flexible approach, allowing for greater use of advanced methods and better adaptability while respecting the original framework and intentions set by the challenge organisers.
@@ 6 Evaluation
We evaluated a prototype of AccelerQ for its ability to optimise further ADAPT-QSCI and QCELS algorithms by suggesting better hyperparameters tailored per system, in this report focusing on random search without a crossover operator.
6.1 Methodology
We demonstrated our idea by optimising hyperparameters of two quantum eigensolver algorithms. We used 4- to 16-qubit systems to train the models. We deployed the models on 20-, 24- and 28-qubit systems. We aim to understand:
RQ1: To what extent can AccelerQ’s optimisation of hyperparameters alone accelerate and improve the efficiency and accuracy of QE algorithms on NISQ devices in terms of error, and system size?
@@ 6.1 Methodology
In our evaluation, we applied our methodology to 16 larger systems of 20-, 24- and 28-qubit with known lowest energy levels [^6] In the case of the open-source systems, we have only a close estimation using ADAPT-QSCI hyperparameters results as a baseline., using the two QE implementations: ADAPT-QSCI and QCELS (Section 4) to answer RQ1 and RQ2.
We constructed two models—one for each implementation—using data extracted classically (Section 4) from up to 16-qubit systems. The data extraction and model training were general. The resultant model aimed to predict the optimal hyperparameters for its QE implementation and a system (Hamiltonian). The optimisation process took an implementation, its trained model and a system (a problem to solve) as input, and returned the predicted optimal hyperparameters for this setup.
@@
Our evaluation baseline is the results obtained with an implementation’s default hyperparameters, fixed across all systems in the evaluation. The hyperparameters were specific per implementation, were optimised, and their performance was compared against the baseline. Apart from the hyperparameters, ADAPT-QSCI and QCELS implementations were provided with ham and number_qubits, the input system and the required number of qubits for the corresponding Hamiltonian, and the flag is_classical being set to True during the data collection phase, and either True or False otherwise (exact classical simulation or MPS). We did not optimise this part.
@@
ADAPT-QSCI. The ADAPT-QSCI implementation [49] includes several hyperparameters that control its operation. These hyperparameters have default values, which we used as our baseline for comparison. num_pickup (default: 100) and coeff_cutoff (default: 0.001) parameters control the terms retained or removed from the Hamiltonian compressed representation. self_selection indicates whether self-selection is enabled (default: False), which forces the algorithm to work in the subspace with the correct Hamming weight. iter_max is the maximum total number of iterations (default: 100). sampling_shots is the number of sampling shots for measurements per iteration (default: $10^{5}$). atol is the absolute tolerance for convergence criteria (default: $61e-$ 6). final_sampling_shots_coeff is the number of shots for the calculations if the same operator appears twice or the operator parameter is close to zero (default: 5). num_precise_gradient is the number of operators from the pool to calculate the gradient with the full Hamiltonian, after an initial approximate calculation is carried out with the truncated Hamiltonian (default: 128). reset_ignored_inx_mode specifies after how many iterations previously used operators can be reused (default: 0). These default values are the baseline for configuring the ADAPT-QSCI implementation, but these can be overridden (e.g.) when using AccelerQ suggestions tailored per system.
@@
QCELS. The QCELS hyperparameters are a different set of arguments. This includes the following: ham_terms is the number of individual terms that are retained in the Hamiltonian after truncation. The Hamiltonian is transformed to a qubit Hamiltonian formed from a linear combination of Pauli strings, the truncation then retains the largest ham_terms of these and the rest are discarded. The default value is set to be 200 and we search for its optimal value between random.randint(50, 1000). ham_cutoff is the same as coeff_cutoff in ADAPT-QSCI implementation, setting a minimum value for the retained coefficients, discarding all terms in the Hamiltonian with coefficients lower than this threshold. delta_t is the time step for the simulation or evolution of the system (the default is 0.03; delta_t = random.uniform(1e-3, 0.3)), the expectation value of the time evolution operator is calculated at a set of times starting from zero and separated by this value. n_Z is the total number of points used in fitting the time evolution and so n_Z11-1- 1 is the total number of points at which the expectation value is evaluated. The default value is 10; n_Z = random.randint(5, 25). alpha is a scalar parameter that determines how much smaller the parameters introduced in each step of the fitting procedure are than those introduced in the previous step e.g. $r_{2}^{(2)}$, $\theta_{2}^{(2)}$ and $r_{2}^{(3)}$, $\theta_{2}^{(3)}$, are bounded by $\alpha r_{1}^{(1)},\pm\alpha\theta_{1}^{(1)}$ (default is $0.8$; alpha = random.uniform(0.5, 1)).
@@ 6.3 Results
Table 1: Predicted optimal hyperparameters (A-J col.) with the first row stating the default values for ADAPT-QSCI implementation (A: num_pickup, B: coeff_cutoff, C: self_selection, D: iter_max, E: sampling_shots, F: atol, G: final_sampling_shots_coeff, H: num_precise_gradient, and I: reset_ignored_inx_mode).
System A B C D E F G H I
Default 100 0.001 0 100 100000 1.00E-06 5 128 0
20qubits_00 985 7.64647e-03 0 344658 49458 5.67168e-05 3 77 1
20qubits_01 807 9.41494e-03 1 382109 964853 9.00118e-05 6 79 64
20qubits_02 934 4.56117e-03 0 478268 875460 4.27314e-05 8 294 83
20qubits_03 551 1.13139e-03 1 802530 106374 7.23056e-05 1 161 79
20qubits_04 182 4.06265e-03 0 398085 61950 7.69414e-05 7 194 11
20qubits_05 593 8.46933e-04 1 278431 148423 6.66628e-05 7 158 58
24qubits_05 67 7.77741e-03 1 655715 676443 3.37411e-05 8 176 89
24qubits_06 292 4.00805e-03 0 762863 603105 6.33079e-05 3 159 82
24qubits_07 80 4.06535e-03 0 659716 711486 9.45728e-05 4 37 0
24qubits_08 215 6.59156e-03 1 885413 291401 1.74522e-05 1 247 81
24qubits_09 716 7.29064e-03 0 401350 929361 6.82934e-05 5 257 65
28qubits_00 197 1.42217e-03 1 460054 209518 7.62522e-05 7 287 54
28qubits_01 1000 6.58567e-03 0 615787 300938 1.04770e-05 8 224 98
28qubits_02 531 5.88032e-03 1 402352 258683 4.81220e-06 8 277 47
28qubits_03 629 4.35568e-03 0 429007 82966 3.89304e-05 8 239 8
28qubits_04 698 7.96341e-03 0 221061 633412 7.86532e-05 4 68 89
@@
Table 2: Predicted optimal hyperparameters (A-E cols.): the approximate size of the Hamiltonian as a vector of terms, A: delta_t, B: n_Z, C: ham_terms, D: ham_cutoff, and E: alpha).
System A B C D E
Default 0.03 10 200 1e-9 0.8
20qubits_00 1.37929e-01 14 877 6.77275e-03 9.76144e-01
20qubits_01 2.87833e-01 8 69 9.23310e-04 5.26429e-01
20qubits_02 1.05366e-01 10 989 4.31098e-04 8.77018e-01
20qubits_03 1.04235e-01 16 68 6.91101e-03 5.12876e-01
20qubits_04 1.31645e-01 19 971 2.10235e-03 7.76523e-01
20qubits_05 1.98379e-01 7 187 8.34526e-03 5.33536e-01
24qubits_05 2.08699e-01 24 933 5.75035e-03 6.20185e-01
24qubits_06 1.75486e-01 10 398 7.28888e-03 9.72887e-01
24qubits_07 1.80166e-01 11 286 1.06348e-03 9.75924e-01
24qubits_08 2.72722e-01 22 453 6.67156e-03 8.29813e-01
24qubits_09 2.09069e-01 24 650 3.99377e-03 5.20032e-01
28qubits_00 7.00174e-03 7 310 6.71154e-03 6.32763e-01
28qubits_01 1.99713e-01 23 309 8.39957e-03 5.54053e-01
28qubits_02 5.46760e-02 12 997 1.22706e-03 5.60243e-01
28qubits_03 4.23353e-02 6 553 8.40543e-03 8.66162e-01
28qubits_04 2.33005e-01 25 280 1.30071e-04 5.98526e-01
Table 3: Hamiltonian sizes of seven Hamiltonian representations. The Hamiltonian sizes in the five rightmost columns are post-Jordan-Wigner Transformation. We marked the open-source molecular Hamiltonians with an asterisk. (1) FermionOperator term count, in the original Hamiltonian representation (#terms col.); (2) The size of a vector of floating point numbers in NumPy of the compressed, flattened FermionOperator representation for machine learning processing (#floats in ML model col.); (3) initial Hamiltonian term count (Init. col.), and its size compressed via cutoff and term selection of (4) default and (5) optimised ADAPT-QSCI, and (6) default and (7) optimised QCELS hyperparameters setup (Round ADAPT-QSCI, Round opt. ADAPT-QSCI, Truncated QCELS, and Truncated opt. QCELS columns).
#terms in Hamiltonian (Jordan-Wigner)
System #terms
#floats in
ML model
Init.
Round
ADAPT-QSCI
Round, Opt.,
ADAPT-QSCI
Truncated
QCELS
Truncated,
Opt.,QCELS
20qubits_00 63636 91 14537 100 985 200 877
20qubits_01 54717 287 14537 100 807 200 69
20qubits_02 54723 445 14535 100 934 200 989
20qubits_03 54710 338 14551 100 551 200 68
20qubits_04 63636 65 14537 100 182 200 971
20qubits_05(*) 46 46 67 67 67 66 66
24qubits_05(*) 56 48 81 81 67 80 80
24qubits_06(*) 56 46 81 81 81 80 80
24qubits_07(*) 56 50 81 81 80 80 80
24qubits_08(*) 56 51 81 81 81 80 80
24qubits_09(*) 56 47 81 81 81 80 80
28qubits_00 98700 167 42982 100 197 200 310
28qubits_01 98700 158 42982 100 1000 200 309
28qubits_02 98700 98 42982 100 531 200 997
28qubits_03 98700 144 42982 100 629 200 553
28qubits_04 98700 213 42982 100 698 200 280
@@
Table 1 and Table 2 show the optimal hyperparameters found by AccelerQ with each Hamiltonian (System col.) employing the two models above. The first row shows the default values of the implementation. Table 1 shows the ADAPT-QSCI’s optimal hyperparameters, with col. A-I represents the predicted optimal values of the hyperparameters. Similarly, Table 2 presents the optimal values for the hyperparameters of the QCELS implementation in col. A-E. The value presented in Table 1 for the predicted iter_max was capped during execution [^8] The platform automatically stopped the computation once the maximum number of shots, 10 000 0001000000010\,000\,00010 000 000, was reached. to ensure that we do not have unlimited resources. In practice, the iter_max has the same limit as with the default parameters (i.e. iter_max=1e7/sampling_shots using the values in col. E and col. D).